OpenStudy (unklerhaukus):

@mathematics What do you get when you add 3 apples to 6 oranges?

OpenStudy (anonymous):

You have 0 fruits. Because I will eat them all! Jk, you get 3+6 =9 fruits...? haha.

OpenStudy (unklerhaukus):

well if you have eaten all my apples and oranges you'll get indigestion

OpenStudy (anonymous):

:D worth it! delicious. i am eating an apple right now.

OpenStudy (unklerhaukus):

3a+6o= zero (cause someone ate them all)

OpenStudy (anonymous):

3a + 6o + amanda = nom + time = indigestion